OpenStudy (anonymous):

Simplify Squareroot 5/8

OpenStudy (anonymous):

like this \[\sqrt{\frac{ 5 }{ 8 }}\]?

OpenStudy (anonymous):

2 * 4 =8.... what is the square root of 4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2

OpenStudy (anonymous):

well Is this the same thing than\[\sqrt{\frac{ 5 }{ 2 \times 4 }}\]

OpenStudy (anonymous):

i guess yes

OpenStudy (anonymous):

so if you said the square root of 4 was 2 could you take the 2 out than like this \[\frac{ 1 }{ 2 }\sqrt{\frac{ 5 }{ 2 }}\]

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

all I did was take the square root of 4 in the denominator which is 1/2 ...

OpenStudy (anonymous):

is that the answer than?

OpenStudy (anonymous):

what is the answer the ... one I wrote above.... Do you think it is ?

OpenStudy (anonymous):

\[4\sqrt10\] thats one of the answers

OpenStudy (anonymous):